OpenStudy (anonymous):

10 - 2x = 4x + 34 answer please

jhonyy9 (jhonyy9):

-6x=24 x=24/(-6)=-4

OpenStudy (lilg132):

10 - 2x = 4x + 34 -2x - 4x = 34 - 10 -6x = 24 -6x / 6 = 24 / 6 -x = 4 swap them around to get x as a positive -4 = x